# IFT-UAM/CSIC-24-116

### v1

@@
1Instituto de Física Teórica, IFT-UAM/CSIC, 28049 Madrid, Spain
2Departamento de Física Teórica, Universidad Autónoma de Madrid, 28049 Madrid, Spain
@@ 1 Introduction
In this article, we develop a modular procedure for estimating the DM model parameters from DM direct detection data. The algorithm is built upon Simulation Based Inference (SBI) techniques, more specifically, the so called Truncated Marginal Neural Ratio Estimation (TMNRE) [38] method. These methods take advantage of machine learning algorithms  [^1] Other machine learning techniques applied to DM direct detection can be found in Refs. [39, 40, 41, 42] for phenomenological studies, and in Refs. [43, 44, 45, 46, 47, 48] for experimental analyses. to perform a Bayesian estimation of model parameters without the need to assume a given likelihood. This method has several advantages. First, once the model is trained, it can be applied to new data to obtain the posterior probabilities much faster than traditional methods. Second, the marginal distributions are computed automatically without the need of integrating the full posterior, which considerably speeds up the computation.
@@ 2 Truncated Marginal Neural Ratio Estimation for Parameter Reconstruction
Machine learning methods have become a very useful tool to analyse large datasets, aiming to extract as much information as possible from the data itself. In the last years, several techniques have been developed that improve Bayesian inference by analysing the simulations computed with the forward model $\mathcal{F}({\bf\Theta})$ [38, 58, 59, 60]. Among the different SBI methods, there are some which are specifically designed to directly estimate the posterior function [59], the likelihood function [60] or the likelihood to evidence ratio as presented in Ref.[38]. In this work, we make use of the TMNRE method [38], which uses neural networks to estimate the marginal likelihood-to-evidence ratio, $r({\bf X},{\bf\Theta})$, defined as
@@ 3.2 Preparation of the datasets
In order to simulate different interaction types in a direct detection experiment, we modified a software developed by the XENON Collaboration [68, 69, 70], to match XENONnT specifications [51, 8] (size of the TPC, fiducial mass, drift field, electron lifetime, photon detection probability $g_{1}$, effective charge gain $g_{2}$, background levels, and a projected exposure of 20 ton yr). To model the backgrounds mentioned above, we used the templates included in the package, except the solar neutrino contribution that we replaced with SNuDD [71], a Python package with accurate computations of solar neutrino spectra and scattering rates at direct detection (DD) experiments in the presence, or not, of non-standard neutrino interactions. Notice that these background templates are not the ones that the XENON collaboration employs for their analysis, however they are very useful to probe the robustness and reach of the proposed method in the current manuscript. A more detailed description of the backgrounds or replacing the detector response/simulator can be easily implemented, however we do not expect significant changes in our main conclusions.
In order to generate the DM signal, we must provide the DM differential rate to the XENON simulator. To do this, we used WimPyDD [67], a code that calculates accurate predictions of nuclear scattering expected rates and spectra for direct detection experiments, for any given NR-EFT operator and parameters, $m_{\chi}$, $\sigma$ and $\theta$.
@@ 3.3 Training of the machine learning algorithms
As is common in machine learning techniques, in order to achieve a better performance all the data was normalised between 0 and 1. In addition, as explained in Ref. [65], when dealing with complex (or high-dimensional) data it is beneficial to perform a data compression before the actual ratio estimation. In our case, for the differential rate and cS1–cS2 representations, we added an extra neural network (a dense neural network for the differential rate analysis and a convolutional neural network for the cS1–cS2 case) to reduce the dimensionality of the problem to 5 and 10 parameters, respectively. We have checked that increasing the final number of parameters does not improve the performance. Also, we have found that the final results are robust to the choice of architectures of the compression networks.
@@ 4 Validation and reconstruction of DM parameters
The analyses are consistent with the true values of the benchmark point (shown as a black and yellow diamond). As expected, the addition of new data is crucial to put better constraints on the different parameters. For example, analysing only the total rate it is not enough for determining the mass of such hypothetical dark matter particle (uppermost panel) because for every mass value considered a very similar total number of events can be achieved. This picture changes drastically when including the differential rate and the full cS1–cS2 information, allowing not only to establish a lower bound for the mass but to delimit the cross section. However, the degeneracy in $\theta$ generally can not be resolved for any benchmark point with a single target. Including information from other experiments with complementary target nuclei, i.e., with different suppression phases (for xenon $\theta\simeq 0.17$), would help constraint this parameter.
Figure 2: Posterior distributions for a dark matter particle with an $\mathcal{O}_{1}$ interaction, $m_{\chi}=50$ GeV, $\sigma^{SI}=2\times 10^{-47}$ cm2, and $\theta=\pi/4$. In the diagonal we present the $1$D mass, cross-section, and phase marginal posterior distributions, while in the lower left panels we show the $2$D marginal posterior distributions. In orange are presented the results obtained analysing the total number rate, in blue the results when adding also the differential rate information and in green the results when adding also the full cS1–cS2 expected signal. The grey areas correspond to the posteriors from an MCMC analysis.
@@
Another crucial characteristic to compare is the computational cost of both approaches. For each benchmark point considered in this work the traditional MCMC analysis took $\sim 2$ days of computation in a normal laptop. On the other hand, the TMNRE method only took $\sim 10$ seconds to fully estimate the required posteriors once the algorithm is already trained (the training procedure required $\sim$ 30 minutes and had to be done only once). This computational speed-up opens the window to analyse an otherwise computationally prohibitive number of benchmark points. In addition, it also makes possible to test different data representations and/or experiments in a reasonable amount of time since it is trivial to combine different analyses in the TMNRE context (see Section 2.1). On the contrary, in an MCMC analysis one has to assume a new combined likelihood (that can be the multiplication of the individual likelihoods) and re-run all the chains, which would also take a few days of computation per benchmark point.
Figure 3: Posterior distributions for a DM particle with an $\mathcal{O}_{4}$ interaction with $m_{\chi}=50$ GeV, $\sigma^{SD}=2.3\times 10^{-40}$ cm2, and $\theta=0$. Colours and labels are the same as in Fig. 2.
In Fig. 3 we show the posterior distributions obtained for a DM particle with $\mathcal{O}_{4}$ interaction, $m_{\chi}=50$ GeV, $\sigma^{SD}=2.3\times 10^{-40}$ cm2, and $\theta=0$. As in the $\mathcal{O}_{1}$ case, the TMNRE and MCMC methods provide consistent results, and both recover the true values of the analysed parameters.
@@ 4.1 Reconstruction of benchmark points
Using these criteria, we have studied the regions where the DM parameters can be reconstructed in a future xenon experiment. The results are shown in Figs. 4 and 5 where, using new independent datasets, we have analysed ($m_{\chi}$, $\sigma$) slices with fixed values of $\theta$, for $\mathcal{O}_{1}$ and $\mathcal{O}_{4}$ operators, respectively. We have divided each slice in a $30\times 30$ grid and for each point we generated five synthetic observations with the same parameters (since our simulator includes statistical fluctuations, those five synthetic observations turn out to be different from each other). We have analysed each synthetic observation with the already trained neural network (see Section 3.3) and obtained the $1$D marginal posterior $P^{j}(\Theta_{i}|{\bf X}^{j})$, which we use to calculate $\mathcal{P}_{\Theta_{i}}^{j}$, where $j=1$ to $5$ represents each synthetic observation. Finally, we have applied the mentioned criteria to the mean of synthetic observation results, i.e., $\langle\mathcal{P}_{\Theta_{i}}\rangle=\sum_{j=1}^{5}\mathcal{P}_{\Theta_{i}}^{j}/5>0.9$.
@@
In Figs. 4 and 5 the regions above the solid green lines correspond to points in the parameter space where the scattering cross section ($\sigma^{SI}$ in Fig. 4 or $\sigma^{SD}$ in Fig. 5) can be reconstructed in a future xenon detector. The sensitivity can vary up to two orders of magnitude for different values of $\theta$, due to the presence of accidental cancellations between the proton and neutron contributions (also referred to as blind spots in the literature) in the calculation of the total rate. For the spin-independent operator $\mathcal{O}_{1}$ the cancellation occurs near $\theta\sim 0$, whereas for $\mathcal{O}_{4}$, it is located at $\theta\sim\pi/4$.
@@ A.1 MCMC vs TMNRE
Appendix B CADDENA: A Python implementation.
@@
In the current version, all the data and the neural networks models used in this work are available without the need of re-training. This includes the simulations and the neural networks weights of the models for dark matter with $\mathcal{O}_{1}$ and $\mathcal{O}_{4}$ interactions for the three data representations considered. All these features allows the reproduction of the results presented in this paper. In the near future we plan to incorporate more models to analyse the complete set of NR-EFT operators and the response of other experimental facilities. All the available models can be listed and loaded with the following commands:

### v2

@@
1Instituto de Física Teórica, IFT-UAM/CSIC, Campus de Cantoblanco, 28049 Madrid, Spain
2Departamento de Física Teórica, Universidad Autónoma de Madrid, Campus de Cantoblanco, 28049 Madrid, Spain
@@ 1 Introduction
In this article, we develop a modular procedure for estimating the DM model parameters from DM direct detection data. The algorithm is built upon Simulation Based Inference (SBI) techniques, more specifically, the so called Truncated Marginal Neural Ratio Estimation (TMNRE) [38] method. These methods take advantage of machine learning algorithms  [^1] Other machine learning techniques applied to DM direct detection can be found in Refs. [39, 40, 41, 42, 43] for phenomenological studies, and in Refs. [44, 45, 46, 47, 48, 49] for experimental analyses. to perform a Bayesian estimation of model parameters without the need to assume a given likelihood. This method has several advantages. First, once the model is trained, it can be applied to new data to obtain the posterior probabilities much faster than traditional methods. Second, the marginal distributions are computed automatically without the need of integrating the full posterior, which considerably speeds up the computation.
@@ 2 Truncated Marginal Neural Ratio Estimation for Parameter Reconstruction
Machine learning methods have become a very useful tool to analyse large datasets, aiming to extract as much information as possible from the data itself. In the last years, several techniques have been developed that improve Bayesian inference by analysing the simulations computed with the forward model $\mathcal{F}({\bf\Theta})$ [38, 59, 60, 61, 43]. Among the different SBI methods, there are some which are specifically designed to directly estimate the posterior function [60], the likelihood function [61] or the likelihood to evidence ratio as presented in Ref.[38]. In this work, we make use of the TMNRE method [38], which uses neural networks to estimate the marginal likelihood-to-evidence ratio, $r({\bf X},{\bf\Theta})$, defined as
@@ 3.2 Preparation of the datasets
In order to simulate different interaction types in a direct detection experiment, we modified a software originally created to simulate XENON1T data, developed by the XENON Collaboration [69, 70, 71], to match XENONnT specifications [52, 8] (size of the TPC, fiducial mass, drift field, electron lifetime, photon detection probability $g_{1}$, effective charge gain $g_{2}$, background levels, and a projected exposure of 20 ton yr). To model the backgrounds mentioned above, we used the templates included in the package, except the solar neutrino contribution that we replaced with SNuDD [72], a Python package with accurate computations of solar neutrino spectra and scattering rates at direct detection (DD) experiments in the presence, or not, of non-standard neutrino interactions. Notice that these background templates are not the ones that the XENON collaboration employs for their analysis, however they are very useful to probe the robustness and reach of the proposed method in the current manuscript. This software is one of the few public tools that allow to obtain direct detection raw data. We have validated the introduced modifications by comparing the expected number of background events per channel, the expected number of signal events for some DM models, and the exclusion curves with XENON publications [52, 8] by performing a frequentist analysis. It is important to highlight that the method proposed in this work is independent of the simulator or particular direct detection experiment. A more detailed description of the backgrounds or replacing the detector response/simulator can be easily implemented, however, we do not expect significant changes in our main conclusions.
In order to generate the DM signal, we must provide the theoretical DM differential rate to the XENON simulator, i.e. the expected number of events per recoil energy. To do this, we used WimPyDD [68], a code that calculates accurate predictions of nuclear scattering expected rates and spectra for direct detection experiments, for any given target material, NR-EFT operator and parameters, $m_{\chi}$, $\sigma$ and $\theta$. The transferred energy is selected following the expected energy distribution probability, while the propagation and final simulated signal also take into account propagation and efficiency effects (in electron propagation in the xenon medium, extraction, PMT response, light collection), resulting in the stochasticity of the final data.
@@ 3.3 Training of the machine learning algorithms
As is common in machine learning techniques, in order to achieve a better performance all the data was normalised between 0 and 1. In addition, as explained in Ref. [66], when dealing with complex (or high-dimensional) data it is beneficial to perform a data compression before the actual ratio estimation. In our case, for the differential rate and cS1–cS2 representations, we added an extra neural network to reduce the dimensionality of the problem. For the differential rate analysis, we add a dense neural network composed of 5 hidden layers of 500, 1000, 500, 50 and 5 neurons with a ReLu activation function. For the cS1–cS2 case, we used a CNN composed of a 2D convolutional block with 10 $5\times 5$ filters followed by a max pooling with 2 and ReLu activation function, followed by a second convolutional block composed of 20 $5\times 5$ filters, again followed by a max pooling layer. Finally, we add a flatten layer followed by 3 dense hidden layers of 80, 50 and 10 units respectively. These networks reduce the input parameters to 5 and 10, and we have checked that increasing the final number of parameters does not improve the performance. Also, we have analysed how the validation loss after training varies with the number of training realizations. We found that the performance already stabilize when training with around $\sim 15$k realizations. Furthermore, we have found that the final results are robust to the choice of architectures of the compression networks.
@@ 4 Validation and reconstruction of DM parameters
The analyses are consistent with the true values of the benchmark point (shown as a black and yellow diamond). As expected, the addition of new data is crucial to put better constraints on the different parameters. For example, analysing only the total rate it is not enough for determining the mass of such hypothetical DM particle (uppermost panel) because for every mass value considered a very similar total number of events can be achieved. This picture changes drastically when including the differential rate and the full cS1–cS2 information, allowing not only to establish a lower bound for the mass but to delimit the cross section. However, the degeneracy in $\theta$ generally can not be resolved for any benchmark point with a single target. Including information from other experiments with complementary target nuclei, i.e., with different suppression phases (for xenon $\theta\simeq 0.17$), would help constraint this parameter.
Figure 2: $90\%$ probability regions of the posterior distributions for a dark matter particle with an $\mathcal{O}_{1}$ interaction, $m_{\chi}=50$ GeV, $\sigma^{SI}=2\times 10^{-47}$ cm2, and $\theta=\pi/4$. In the diagonal we present the $1$D mass, cross-section, and phase marginal posterior distributions, while in the lower left panels we show the $2$D marginal posterior distributions. In orange are presented the results obtained analysing the total number rate, in blue the results when adding also the differential rate information and in green the results when adding also the full cS1–cS2 expected signal. The grey areas correspond to the posteriors from an MCMC analysis.
@@
Another crucial characteristic to compare is the computational cost of both approaches. For each benchmark point considered in this work the traditional MCMC analysis took $\sim 2$ days of computation in a normal laptop. On the other hand, while the TMNRE method also needs $\sim 2$ days to generate the required dataset to train the ML model, it only takes $\sim 10$ seconds to fully estimate the required posteriors once the algorithm is already trained (the training procedure required $\sim$ 30 minutes and had to be done only once). This computational speed-up opens the window to analyse an otherwise computationally prohibitive number of benchmark points. In addition, it also makes possible to test different data representations and/or experiments in a reasonable amount of time since it is trivial to combine different analyses in the TMNRE context (see Section 2.1). On the contrary, in an MCMC analysis one has to assume a new combined likelihood (that can be the multiplication of the individual likelihoods) and re-run all the chains, which would also take a few days of computation per benchmark point.
Figure 3: Empirical coverage of the posterior distributions as a function of the nominal credibility region for different values of $z_{p}=1-\alpha$. Left, middle and right columns correspond to $m_{\chi}$, $\sigma^{\textit{SI}}$ and $\theta$ respectively. Upper, middle and lower rows correspond to total rate, differential rate and cS1-cS2 respectively.
As an additional consistency check, we compute the empirical coverage of the posterior distributions. Although there are alternative ways of computing the coverages [77], we follow the procedure described in [66] that is already implemented in the SWYFT package. The idea behind this procedure is that, when randomly drawing observations from ${\bf X},{\bf\Theta}\sim p({\bf X}|{\bf\Theta})p({\bf\Theta})$, it is expected that the true parameters ${\bf\Theta}$ will fall outside the $1-\alpha$ confidence region only in the $\alpha$of the cases. Any departure from this implies that the posterior distributions are over-confident (its empirical coverage is lower than the nominal credibility, for example if a 95% region is constructed, but in simulations it only covers the true parameter 90% of the time) or conservative (its empirical coverage is higher than the nominal credibility, for example if the 95% region actually covers the true parameter 98% of the time). As estimating the posterior distributions may be very expensive with traditional methods, these kinds of tests are usually impossible to do. In contrast, once trained, the TMNRE method allows for very fast estimation of posterior distributions, enabling further coverage checks. In Fig. 3, we show the empirical coverage as a function of the nominal credibility region for different values of $z_{p}=1-\alpha$. It can be seen that the estimated posteriors of all the parameters (left, middle, and right columns correspond to $m_{\chi}$, $\sigma^{\textit{SI}}$ and $\theta$, respectively) and all the data representations (upper, middle, and lower rows correspond to total rate, differential rate and cS1-cS2, respectively) are well calibrated. The only noticeable exceptions are an over-confident region (for $z_{p}\gtrapprox 3$) in the $\sigma^{\textit{SI}}$ posterior when using the differential rate, and a conservative region ($0.5\gtrapprox z_{p}\gtrapprox 2.5$) in the $\sigma^{\textit{SI}}$ posterior when using cS1-cS2 data.
Figure 4: $90\%$ probability regions of the posterior distributions for a DM particle with an $\mathcal{O}_{4}$ interaction with $m_{\chi}=50$ GeV, $\sigma^{SD}=2.3\times 10^{-40}$ cm2, and $\theta=0$. Colours and labels are the same as in Fig. 2.
In Fig. 4 we show the $90\%$ probability regions obtained for a DM particle with $\mathcal{O}_{4}$ interaction, $m_{\chi}=50$ GeV, $\sigma^{SD}=2.3\times 10^{-40}$ cm2, and $\theta=0$. As in the $\mathcal{O}_{1}$ case, the TMNRE and MCMC methods provide consistent results, and both recover the true values of the analysed parameters.
@@ 4.1 Reconstruction of benchmark points
Using these criteria, we have studied the regions where the DM parameters can be reconstructed in a future xenon experiment. The results are shown in Figs. 5 and 6 where, using new independent datasets, we have analysed ($m_{\chi}$, $\sigma$) slices with fixed values of $\theta$, for $\mathcal{O}_{1}$ and $\mathcal{O}_{4}$ operators, respectively. We have divided each slice in a $30\times 30$ grid and for each point we generated five synthetic observations with the same parameters (since our simulator includes statistical fluctuations, those five synthetic observations turn out to be different from each other). We have analysed each synthetic observation with the already trained neural network (see Section 3.3) and obtained the $1$D marginal posterior $P^{j}(\Theta_{i}|{\bf X}^{j})$, which we use to calculate $\mathcal{P}_{\Theta_{i}}^{j}$, where $j=1$ to $5$ represents each synthetic observation. Finally, we have applied the mentioned criteria to the mean of synthetic observation results, i.e., $\langle\mathcal{P}_{\Theta_{i}}\rangle=\sum_{j=1}^{5}\mathcal{P}_{\Theta_{i}}^{j}/5>0.9$. It is important to clarify that each of the five synthetic observations corresponds to the expected real data if DM has the specific parameters under analysis. In this way, the reconstruction regions take into account possible statistical fluctuations of the data.
@@
In Figs. 5 and 6 the regions above the solid green lines correspond to points in the parameter space where the scattering cross section ($\sigma^{SI}$ in Fig. 5 or $\sigma^{SD}$ in Fig. 6) can be reconstructed in a future xenon detector. The sensitivity can vary up to two orders of magnitude for different values of $\theta$, due to the presence of accidental cancellations between the proton and neutron contributions (also referred to as blind spots in the literature) in the calculation of the total rate. For the spin-independent operator $\mathcal{O}_{1}$ the cancellation occurs near $\theta\sim 0$, whereas for $\mathcal{O}_{4}$, it is located at $\theta\sim\pi/4$.
@@
The selection of the integration limits of Eq. 10 is not unique. This is specially relevant in the DM mass case, for which there is no zero-signal limit (contrarywise, for the cross-section analysis, the zero-signal case can be obtained for small enough $\sigma$). However, this analysis provides an initial insight into the regions of the parameter space where these properties could be measured. In order to test the impact of varying the integration limits, we performed the same analysis varying the limits in the ranges $\sigma^{SI}_{min}\in[10^{-49.3},10^{-48.7}]$ cm2, $\sigma^{SD}_{min}\in[10^{-42.3},10^{-41.7}]$ cm2, and $m_{\chi,max}\in[400,650]$ GeV. When further increasing $m_{\chi,max}$, we approach to the borders of our training set and therefore it is equivalent to the $m_{\chi,max}=\infty$ case (dashed green lines). The results obtained when varying the integration limits are displayed as green shaded bands around the corresponding lines on Figs. 5 and 6.
@@ A.1 MCMC vs TMNRE
Appendix B CADDENA: A Python implementation
@@
In the current version, all the data and the neural networks models used in this work are available without the need of re-training. This includes the simulations and the neural networks weights of the models for DM with $\mathcal{O}_{1}$ and $\mathcal{O}_{4}$ interactions for the three data representations considered. All these features allows the reproduction of the results presented in this paper. In the near future we plan to incorporate more models to analyse the complete set of NR-EFT operators and the response of other experimental facilities. All the available models can be listed and loaded with the following commands: